OpenStudy (anonymous):

Hello. Can I start my iteration when n=-1, no? Let's take for an example this series of expansion: 1/x +1 + x/2! +x^2/3! + ... + x^n/n! Can I say "the algebraic sum from n=-1 to n=infinity"?

OpenStudy (anonymous):

I am asking this because my instructor marked my answer wrong and he won't talk to me.

OpenStudy (anonymous):

No, the index n can only be positive. And what does (-1)! mean?

OpenStudy (anonymous):

that's why i put (n+1)! haha

OpenStudy (anonymous):

Ahhhhhh...gotcha. I like the idea. But you have to have positive indices.

OpenStudy (anonymous):

so should I have just removed my lnx from my series representation?

OpenStudy (anonymous):

No, the teacher was correct. You must write it as an intergral, or a sum using sigma.

OpenStudy (anonymous):

the question is evaluate the integral of e^-x (1/x)dx

OpenStudy (anonymous):

Exactly, you have to leave it as an intergral equal to a sum.

OpenStudy (anonymous):

Can I do that? Oh. I never know. XD

OpenStudy (anonymous):

but can I use lnx + series representation, too?

OpenStudy (anonymous):

Should equal the sum: \[\sum_{n=0}^{\infty}\frac{x^n}{(n!)x}=\sum_{n=0}^{\infty}\frac{x^{n-1}}{n!} \] Which is easy to integrate.

OpenStudy (anonymous):

You can bring the x in the sum because its being summed over n, not x.

OpenStudy (anonymous):

And since you're integrating with respect to x the n! is a "constant" of sorts and can be neglected (there is no antiderivative for n!).

OpenStudy (anonymous):

Or am I inventing math?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+e^(-x)%2Fx Seems plausible.

OpenStudy (anonymous):

(-x)^n / n*n! is this it?

OpenStudy (anonymous):

+C???????

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{(-x)^{n}}{x(n!)}\] You can't bring n's in arbitrarily because its summed over n.

OpenStudy (anonymous):

I got the integral of that (-x)^n ln x / n!

OpenStudy (anonymous):

but whenever i iterate it, all terms contain "lnx" -.-

OpenStudy (anonymous):

Ohhhhh I see what you mean. I don't know how to help that man.

OpenStudy (anonymous):

Oh thanks. my question is answered anyway. :D thanks!